# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ 3.2 Data model, objective, and stochastic-gradient oracle
We report results both in iterations $T$ and in oracle calls $N_{T}$, since the control/OR questions in naturally optimize over batch allocation under a fixed $N_{T}$.
@@ 3.3 Likelihood losses: Fisher information and the induced metric
When $\ell(\theta;z)=-\log p_{\theta}(z)$ for a parametric family $\{P_{\theta}:\theta\in\Theta\}$ with density $p_{\theta}$, the score is $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ and the loss gradient is the negative score, $\nabla_{\theta}\ell(\theta;z)=-s_{\theta}(z)$. The intrinsic noise geometry reduces to Fisher information.
@@
Differentiate $\int p_{\theta}(z)\,dz=1$ to obtain $0=\int\nabla_{\theta}p_{\theta}(z)\,dz=\int p_{\theta}(z)s_{\theta}(z)\,dz$. For (ii), differentiate $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$ and use the identity $\nabla_{\theta}\mathbb{E}_{\theta}[f(Z)]=\mathbb{E}_{\theta}[\nabla_{\theta}f(Z)+f(Z)s_{\theta}(Z)^{\top}]$ with $f=s_{\theta}$. ∎
@@ 3.6 Local regularity conditions
Remark 3.11 (What these assumptions buy us).
Assumption 3.8 ensures uniform Taylor control and moment bounds for the martingale noise. Assumption 3.9 gives a locally stable equilibrium and enables Lyapunov drift arguments. Assumption 3.10 ensures the diffusion coefficient is nondegenerate on identifiable directions and that Fisher/Godambe norms yield meaningful local rates.
@@ 4 Fisher/Godambe Alignment of Mini-Batch Noise
Mini-batch gradient noise is not an arbitrary covariance: it is pinned down by (i) the sampling design used to form mini-batches and (ii) the intrinsic covariance of the per-sample estimating equation $\psi(\theta;Z)$ (Fisher in the well-specified likelihood case, Godambe noise more generally). We make this precise in two steps:
@@ 4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
Remark 4.5 (Independence structure and finite-dataset reuse).
Three conditioning levels should be distinguished: (i) conditional on the dataset $z_{1:n}$: mini-batch covariance is $\frac{1}{b}(1-\frac{b}{n})\,S_{n}(\theta)$ (Lemma 4.1), with the finite-population correction (1b/n)(1-b/n) reflecting sampling without replacement from a fixed pool; (ii) conditional on the directing measure $\mu$: mini-batch covariance is $\frac{1}{b}\,G_{\mu}(\theta)$ (Theorem 4.3), with no correction, because the dataset itself is random; (iii) population level: the unconditional covariance integrates over $\mu$ and dataset randomness, and equals $\frac{1}{b}\,G(\theta)$ under identifiability. The alignment result (23) is iteration-wise under fresh sampling: at each iteration $t$, the mini-batch $B_{t}$ is drawn independently of previous batches conditional on $\mu$. When a finite dataset is reused across iterations (i.e., without-replacement sampling from the same $z_{1:n}$ at every step), the conditional covariance acquires the standard finite-population correction (1b/n)(1-b/n). This correction is negligible whenever $b=o(n)$, which is the typical stochastic-approximation regime. It is the population covariance $\frac{1}{b}\,G_{\mu}(\theta)$ that enters the diffusion approximation (Section 5) and all subsequent rate and complexity results.
@@
Equation (23) implies that increasing the batch size $b$ rescales the effective temperature $\tau=\eta/b$ (Section 5) but does not alter the noise ellipsoid: its shape is entirely determined by the statistical geometry $G_{\mu}(\theta)$. This decoupling—scale from the sampling budget, shape from the loss—is the operational insight that drives all subsequent rate and complexity results.
@@
Implication for later sections.
The analysis in Section 5 depends on a single intrinsic matrix $G^{\star}(\theta^{\star})$ (which equals $F^{\star}(\theta^{\star})$ in the well-specified likelihood case; see Remark 3.6). Once this matrix is fixed, the OU diffusion coefficient, Lyapunov balance, and Fisher-metric rates follow mechanically.
@@ 5 Continuous-Time Diffusion Approximation and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion term is not arbitrary: its covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4. Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance. Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation. The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength. Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
@@ 5.2 A diffusion approximation (Euler–Maruyama viewpoint)
Assumption 5.2 is implied by the following checkable conditions: (a) $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b) with $\delta=2$); (b) $G^{\star}(\theta)$ is continuous on $\mathcal{U}$; and (c) the mini-batch is drawn by simple random sampling without replacement from an exchangeable sequence. Under (a)–(c), the conditional Lindeberg condition follows from Lyapunov’s criterion applied to the triangular array of centered per-sample gradients, and the covariance identity follows from Theorem 4.3.
@@
where $W_{s}$ is standard $d$-dimensional Brownian motion. In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$. In the likelihood / correct-specification case, $G^{\star}(\theta)=F^{\star}(\theta)$, so the diffusion covariance is Fisher-structured.
@@
Step 4 (Lindeberg condition). Assumption 3.8(b) gives $\mathbb{E}[\|\zeta_{k+1}\|^{4}\mid\mathcal{F}_{k}]\leq Cm_{4}/b^{2}$. Markov’s inequality then yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition.
Step 5 (tightness). Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion: drift increments are Lipschitz-controlled and $\mathbb{E}[\|\bar{\theta}^{\eta}(t+\delta)-\bar{\theta}^{\eta}(t)\|^{2}]\leq C\delta$.
Step 6 (martingale FCLT). By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$. Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$. Combining drift and martingale convergence identifies the limit as the unique solution of (28). ∎
Remark 5.5.
The SDE (28) should be read as an Euler–Maruyama limit: SGD is an Euler discretization of a diffusion whose covariance is proportional to the intrinsic gradient-noise geometry $G^{\star}(\theta)$, with overall scale set by the temperature $\tau=\eta/b$.
@@ 5.3 The OU regime near a nondegenerate critical point
Lyapunov Balance Law (structural consequence of sampling alignment). The stationary covariance $\Sigma_{\infty}$ of the OU regime is determined by three objects: the curvature $H^{\star}$ (drift), the intrinsic noise geometry $G^{\star}(\theta^{\star})$ (diffusion), and the effective temperature $\tau=\eta/b$ (scale). The Lyapunov equation $H^{\star}\Sigma_{\infty}+\Sigma_{\infty}H^{\star\top}=\tau\,G^{\star}(\theta^{\star})$ is a structural law, not a modeling assumption: the diffusion matrix on the right-hand side is not postulated but implied by the mini-batch sampling design (Theorem 4.3). Once the loss and the sampling mechanism are fixed, all three ingredients are determined, and so is the stationary covariance.
@@
5.4 Fisher-metric coordinates (whitening)
Remark 5.10 (Whitening by the Fisher metric).
In Fisher-aligned settings, it is often convenient to work in whitened coordinates $x_{s}:=(F^{\star}(\theta^{\star}))^{1/2}\Delta_{s}$. Formally, the OU drift becomes
$dx_{s}=-\Big((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}\Big)x_{s}\,ds+\sqrt{\tau}\,(F^{\star})^{1/2}C^{\star}(\theta^{\star})\,dW_{s},$
so contraction is governed by the spectrum of $H^{\star}$ (equivalently, by the spectrum of the conjugated drift matrix $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$, which is similar to $H^{\star}$).
The diffusion limit as an identification tool.
The diffusion approximation derived above is not an end in itself. Its role is to identify the intrinsic volatility geometry $G^{\star}(\theta^{\star})$ that determines the stationary covariance (Corollary 5.8), the Fisher-risk scaling (Section 6), and the oracle complexity (Section 7). Without the alignment established in Section 4, the diffusion covariance would be a free modeling choice; with it, the covariance is forced by the loss and the sampling design, and every downstream quantity inherits this constraint. Two consequences are verified numerically in Section 8: (i) the stationary Fisher-metric risk stabilises at $\mathrm{Tr}(G^{\star}H^{-1})$, confirming the Lyapunov prediction; and (ii) a trace-matched isotropic surrogate reproduces this scalar but fails to reproduce the off-diagonal covariance structure, confirming that the full geometry—not just the temperature—matters.

### 03 Apr 2026 v2

@@ 3.2 Data model, objective, and stochastic-gradient oracle
We report results both in iterations $T$ and in oracle calls $N_{T}$, since control/OR questions naturally optimize over batch allocation under a fixed $N_{T}$.
@@ 3.3 Likelihood losses: Fisher information and the induced metric
When $\ell(\theta;z)=-\log p_{\theta}(z)$ for a parametric family $\{P_{\theta}:\theta\in\Theta\}$ with density $p_{\theta}$, the score is $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ and the loss gradient is the negative score, $\nabla_{\theta}\ell(\theta;z)=-s_{\theta}(z)$. The intrinsic noise geometry reduces to Fisher information at the correctly specified point $\theta^{\star}$.
@@
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
@@ 3.6 Local regularity conditions
Remark 3.11 (Role of assumptions).
Assumptions 3.83.10 provide Taylor/moment control, a locally stable equilibrium, and nondegeneracy of the diffusion coefficient, respectively.
3.7 Key geometric parameters (preview)
Several quantities derived from the local geometry at $\theta^{\star}$ govern the rate and complexity results of Sections 67. We collect their definitions here so the reader can refer to them before encountering the formal theorems.
Definition 3.12 (Fisher-strong-convexity constant).
Let $H^{\star}:=\nabla^{2}L(\theta^{\star})$ and $F^{\star}:=F(\theta^{\star})$ (or $G^{\star}(\theta^{\star})$ for general losses). Define
$\mu_{F}\;:=\;\lambda_{\min}\!\left(\frac{(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}+(F^{\star})^{-1/2}H^{\star}(F^{\star})^{1/2}}{2}\right).$ (18)
Equivalently, $\mu_{F}$ is the smallest eigenvalue of the symmetrized curvature–information interaction $\mathrm{Sym}((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2})$.
Definition 3.13 (Fisher condition number and effective dimension).
$\kappa_{F}\;:=\;\frac{\lambda_{\max}(F^{\star})}{\lambda_{\min}(F^{\star})},\qquad d_{\mathrm{eff}}(F^{\star})\;:=\;\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}\;\in\;(0,d].$ (19)
The quantity $\kappa_{F}$ measures statistical anisotropy rather than Euclidean ill-conditioning; a problem can be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share eigenstructure. The effective dimension $d_{\mathrm{eff}}$ is the stable rank of $F^{\star}$: it equals $d$ when $F^{\star}$ is isotropic and approaches $1$ when a single direction dominates.
Remark 3.14 (Generalized eigenvalue interpretation).
The constant $\mu_{F}$ is linked to the generalized eigenvalues of the pair $(H^{\star},F^{\star})$. If $H^{\star}$ and $F^{\star}$ commute (i.e., share an eigenbasis $\{v_{j}\}$ with respective eigenvalues $h_{j}$ and $f_{j}$), then $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ has eigenvalues $f_{j}^{1/2}\,h_{j}\,f_{j}^{-1/2}=h_{j}$, so the symmetrization is trivial and $\mu_{F}=\min_{j}h_{j}=\lambda_{\min}(H^{\star})$: the Fisher-strong-convexity constant reduces to the ordinary strong-convexity constant. In the non-commuting case, $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ is non-normal (its eigenvalues are still those of $H^{\star}$, but it is no longer symmetric), and the minimum of its quadratic form—i.e., $\mu_{F}$—can drop strictly below $\lambda_{\min}(H^{\star})$. Geometrically, this occurs when the Fisher metric amplifies directions along which curvature is weak, making the effective contraction rate slower than the Euclidean strong-convexity constant. The product $\kappa_{F}\cdot d_{\mathrm{eff}}$ that enters the oracle complexity (Theorem 7.3) therefore measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale.
Table 2: Key notation. All quantities are evaluated locally at or near $\theta^{\star}$ unless noted otherwise.
Symbol Meaning
$L(\theta)=\mathbb{E}[\ell(\theta;Z)]$ Population risk (loss)
$\psi(\theta;z)=\nabla_{\theta}\ell(\theta;z)$ Per-sample estimating equation (gradient)
$g_{B}(\theta)$ Mini-batch gradient estimator
$G(\theta)=\mathrm{Cov}(\psi(\theta;Z))$ Gradient covariance; $G^{\star}=\Pi G\Pi^{\top}$ projected
$F(\theta)=\mathbb{E}[s_{\theta}s_{\theta}^{\top}]$ Fisher information (likelihood case)
$F^{\star}(\theta^{\star})$ Projected Fisher; equals $G^{\star}(\theta^{\star})$ under correct specification
$H^{\star}=\nabla^{2}L(\theta^{\star})$ Hessian at optimum (drift curvature)
$J_{\mathrm{God}}=H^{\top}G^{-1}H$ Godambe (sandwich) information
$\tau=\eta/b$ Effective temperature
$\mu_{F}$ Fisher-strong-convexity constant (Def. 3.12)
$\kappa_{F}$ Fisher condition number (Def. 3.13)
$d_{\operatorname{eff}}(F^{\star})$ Effective dimension / stable rank of $F^{\star}$
@@ 4 Fisher/Godambe Alignment of Mini-Batch Noise
Mini-batch gradient noise is not an arbitrary covariance: it is determined by (i) the sampling design used to form mini-batches and (ii) the intrinsic covariance of the per-sample estimating equation $\psi(\theta;Z)$ (Fisher in the well-specified likelihood case, Godambe noise more generally). We make this precise in two steps:
@@ 4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
Proposition 4.5 (Finite-population correction under dataset reuse).
Under the notation of Theorem 4.3, three conditioning levels yield distinct covariance expressions:
  1. (i)
    Conditional on the dataset $z_{1:n}$ (fixed pool, sampling without replacement): $\mathrm{Cov}(g_{B}(\theta)\mid z_{1:n})=\frac{1}{b}(1-\frac{b}{n})\,S_{n}(\theta)$ (Lemma 4.1).
  2. (ii)
    Conditional on the directing measure $\mu$ (fresh data per iteration): $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=\frac{1}{b}\,G_{\mu}(\theta)$ (Theorem 4.3), with no finite-population correction.
  3. (iii)
    Population level (unconditional): $\mathrm{Cov}(g_{B}(\theta))=\frac{1}{b}\,\mathbb{E}_{\mu}[G_{\mu}(\theta)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}(\theta)\mid\mu])$. This reduces to $\frac{1}{b}\,G(\theta)$ when the conditional mean $\mathbb{E}[\psi(\theta;Z)\mid\mu]$ does not vary with $\mu$ (e.g., under identifiability with a single data-generating distribution).
When a finite dataset is reused across many iterations (i.e., without-replacement sampling from the same $z_{1:n}$ at every step), the conditional covariance carries the factor (1b/n)(1-b/n), which is non-negligible unless $b=o(n)$. The main diffusion, rate, and complexity results (Sections 57) are population-level or fresh-sampling statements that use level (ii).
Proof.
Level (i) is Lemma 4.1. Level (ii) is Theorem 4.3. Level (iii) follows by integrating level (ii) over $\mu$ under identifiability. The total-covariance decomposition $\mathrm{Cov}(g_{B})=\mathbb{E}_{\mu}[\mathrm{Cov}(g_{B}\mid\mu)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}\mid\mu])$ reconciles levels (ii) and (iii), with the second term contributing when $\mathbb{E}[g_{B}\mid\mu]$ varies with $\mu$. ∎
@@
Equation (22) implies that increasing the batch size $b$ rescales the effective temperature $\tau=\eta/b$ (Section 5) but does not alter the noise ellipsoid: its shape is entirely determined by the statistical geometry $G_{\mu}(\theta)$. This decoupling—scale from the sampling budget, shape from the loss—underlies the subsequent rate and complexity results.
@@ 5 Continuous-Time Diffusion Approximation and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4. Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance. Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation. The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength. Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
@@ 5.2 A diffusion approximation (Euler–Maruyama viewpoint)
Assumption 5.2 holds if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b)), $G^{\star}(\theta)$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
@@
where $W_{s}$ is standard $d$-dimensional Brownian motion. In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$. In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ at $\theta^{\star}$, so the diffusion covariance is locally Fisher-structured near the optimum.
@@
Step 4 (Lindeberg and tightness). The fourth moment bound (Assumption 3.8(b)) yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition. Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion and the Lipschitz-controlled drift.
Step 5 (martingale FCLT). By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$. Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$. Combining drift and martingale convergence identifies the limit as the unique solution of (27). ∎
@@ 5.3 The OU regime near a nondegenerate critical point
5.4 Discrete-time stationary covariance: an exact bridge
The OU viewpoint is not the only consequence of the covariance identification. The same $G^{\star}$ governs the exact linearized discrete-time stationary covariance; the continuous-time Lyapunov equation emerges as the small-stepsize asymptotic reduction. This subsection makes the bridge explicit, so that the paper’s conclusions do not rest on the diffusion approximation alone.
Proposition 5.9 (Discrete-time linearized stationary covariance).
Consider the linearized discrete-time recursion around $\theta^{\star}$ with constant stepsize $\eta$ and batch size $b$:
$\Delta_{t+1}=(I-\eta H^{\star})\,\Delta_{t}+\eta\,\zeta_{t+1},\qquad\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]=\frac{1}{b}\,G^{\star}(\theta^{\star})+r_{t},$ (32)
where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1})$ as in (25). If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), then the iterates are geometrically ergodic and the stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfies the discrete Lyapunov equation
$\Sigma_{\eta}=(I-\eta H^{\star})\,\Sigma_{\eta}\,(I-\eta H^{\star})^{\top}+\frac{\eta^{2}}{b}\,G^{\star}(\theta^{\star}),$ (33)
up to the same local remainder terms already used in the diffusion analysis.
Proof.
From (32), compute $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$, using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$. Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the iteration converges to the unique fixed point (33). The spectral radius condition $\rho(I-\eta H^{\star})<1$ is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$. ∎
Remark 5.10 (Continuous-time reduction and discrete-time corrections).
Expanding (33) and dividing by $\eta$, one recovers $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=\frac{\eta}{b}G^{\star}+\eta H^{\star}\Sigma_{\eta}H^{\star\top}$. Sending $\eta\downarrow 0$ with $\tau=\eta/b$ fixed drops the $O(\eta)$ correction term and yields the continuous Lyapunov equation (30). The correction is governed by $\eta\lambda_{\max}(H^{\star})$: when this product is small the two benchmarks nearly coincide; for moderate stepsizes the discrete equation is the correct comparator (Section 8). The covariance identification $\mathrm{Cov}(\xi_{t})\approx b^{-1}G^{\star}$ is logically prior to both analyses: the discrete Lyapunov equation is its exact linearized consequence, and the OU equation is the small-$\eta$ reduction.